What is the shortest distance between (-2,0) and (2,0) without entering Inside of the circle x^2+y^2=1 ?
1.  2√3
2.  2√3+π/3
3.  √3+2π/3
4.  None of these

Grade:12

1 Answers

Aditya saxena
5 years ago
  1. Draw a diagram of circle with origin as centre and Mark all the points on x and y axis
  2. All points are (1,0) (0,1) (-1,0)(0,-1)
  3. As we know min distance between two points will be a line joining (1,0) (2,0) and (-2,0)
Length of tangent will be \sqrt{}S(-2,0) which will when doubled give the exact length where S= x2+y2=1 
2* \sqrt{}\frac{}{}(-2)+0 -1 = 2\sqrt{}3